OpenStudy (anonymous):

define p~q by the equation p~q=p^2q^3-3q then 2~3

OpenStudy (anonymous):

@phi

OpenStudy (phi):

they are saying if you see 2~3, then match that up with p~q in other words, p=2 and q=3 then use those numbers instead of the letters in the formula p^2q^3-3q can you do that?

OpenStudy (anonymous):

yeah okay, I just didnt know what they were asking thnak you

OpenStudy (anonymous):

4(9)-9=27

OpenStudy (phi):

almost. but it is q^3 = 3^3 = 3*3*3

OpenStudy (anonymous):

4(27)-9=99

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you

OpenStudy (phi):

yw